Determine the equation of the line in slope-intercept form which passes through the point:
(11,19) and is parallel to the line joining the points:
(-4, 4) and (-9, -7).

Answer by stanbon(75887)

Determine the equation of the line in slope-intercept form which passes through the point:
(11,19) and is parallel to the line joining the points:
(-4, 4) and (-9, -7).
----------
slope = (4--7)/(-4--9) = 11/5
-----
Form: y = mx + b
Solve for "b":
19 = (11/5)11 + b
----
95/5 = 121/5 + b
b = -26/5
-------
Equation:
y = (11/5)x - (26/5)
======================
Cheers,
Stan H.
=================